# thanks: The first two authors contributed equally

### v1

@@ II.3 Frequency modulation of the binary
In the previous subsection, we argued that an inspiralling intermediate mass binary black hole (IMBBH) in the vicinity of horizon of supermassive black hole (SMBH) acts as a detector for detecting supertranslation (ST) memory. IMBBH loses energy and angular momentum to GWs which can be observed by Laser Interferometer Space Antenna (LISA). The separation of the binary components $L=2a$ (we call it IMBBH separation) at any instant during inspiral can be related to GW frequency $f$ emitted at that time via Keppler’s law as Peters and Mathews (1963); Peters (1964):
@@
where $a$ is the radius of the binary, $z$ is the cosmological redshift of the IMBBH as well as the SMBH, and $M_{\mathrm{bin}}$ is the source-frame total mass of the IMBBH.
Once the supertranslation ‘hair’ is implanted on the central SMBH, the binary separation in IMBBH changes as a result of ST memory as given by Eq. (11). The sudden change in the IMBBH separation leads to a corresponding sudden change in the emitted GW frequency. This instantaneous frequency jump in the GW signal of the IMBBH should be observable. The detection of the ST memory can be quantified by calculating the mismatch between the original GW signal when there is no ST memory and the signal modulated due to the ST memory.
GWs in general relativity have two polarizations i.e., plus $h_{+}$ and cross $h_{\times}$. For quasicircular binaries, the evolution of these polarizations is determined by the intrinsic source parameters ($\vec{\theta}$) i.e., component masses ($m_{1}$, $m_{2}$) and spins ($\vec{\chi}_{1}$, $\vec{\chi}_{2}$); and the extrinsic source parameters ($\vec{\lambda}_{h}$) i.e., luminosity distance $D_{L}$ to the source, inclination angle $\iota$ of the binary with respect to the observer’s line of sight, time $t_{c}$ and phase $\phi_{c}$ at the coalescence. Note that instead of ($m_{1}$, $m_{2}$), it is more natural to use the chirpmass $\mathcal{M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ and mass ratio ${\bf q}=m_{2}/m_{1}$ [^4] Here q should not be confused with the supertranslation charge $q$ in Eq. (11). in the parameter estimation as the shape of gravitational waveform is more sensitive to the chirpmass Cutler and Flanagan (1994a). An interferometric GW detector, such as LISA, records the time-domain strain $h(t;\vec{\theta},\vec{\lambda})$ due to an incoming GW signal, which can be expressed as a linear combination of the plus ($h_{+}$) and cross ($h_{\times}$) polarizations of GWs weighted by the respective antenna responses ($F_{+},F_{\times}$):
@@
$h_{b}(t+\Delta t;\vec{\theta},\vec{\lambda})=F_{+}(\vec{\lambda}_{F})h_{+}(t+\Delta t;\vec{\theta},\vec{\lambda}_{h})$ (13)
$+F_{\times}(\vec{\lambda}_{F})h_{\times}(t+\Delta t;\vec{\theta},\vec{\lambda}_{h})\,.$
@@
$\hat{\mathcal{A}}=\sqrt{{5\over 24}}\sqrt{{4\over 5}}{\sqrt{\eta}M_{bin}^{5/6}(1+z)^{5/6}\over\pi^{2/3}D_{L}}\,.$ (15)
@@
where $t(f)$ is the time corresponding to GW frequency $f$, $f_{\rm in}$ and $f_{\rm fi}$ are respectively, the GW frequencies at which binary is emitting before and after supertranslation hair is implanted on the central SMBH and ${\rm\mathcal{M}_{c}}=M_{\rm bin}\eta^{3/5}$ is the chirp mass of the binary. Note that in order to simplify that calculations, we assume the frequency jump due to ST memory from $f_{\rm in}$ to $f_{\rm fi}$ to be discrete.
III Bayesian Parameter Inference
@@
[…]
Assuming the detector noise to be stationary and Gaussian, the likelihood can be written as Cutler and Flanagan (1994b):
@@
where $d$ is the strain data of the detector and .|$\langle$.|$\rangle$ . | . ⟩ is the noise-weighted inner product defined as:
[…]
@@
[…]
@@
The lower and upper cutoff frequencies ($f_{\rm low}$ and $f_{\rm high}$) used in calculating the integral in Eq. (21) are determined by the detector sensitivity, observation time, and the source properties. The $f_{\rm low}$ is defined as $f_{\rm low}={\rm max}[10^{-4},f_{\rm year}]$, where $f_{\rm year}$ is the GW frequency ${\rm T_{obs}}$ years before the binary emits at the frequency corresponding to the innermost stable circular orbit (ISCO). This frequency, in the quadrupole radiation approximation, is determined by the chirp mass ${\mathcal{M}_{c}}$ and the observation time prior to ISCO $T_{\rm obs}$ as Berti et al. (2005):
@@
We assume $T_{\rm obs}=4$ year in our calculation. The upper cut-off frequency is given by $f_{\rm high}={\rm min}[0.1,f_{\rm ISCO}]$, where $f_{\rm ISCO}=1/({6^{3/2}\pi(1+z)M_{\rm bin}})$ is the GW frequency corresponding to ISCO of a test particle orbiting around a Schwarzschild black hole of mass $M_{\rm bin}$.
We have displayed in Fig. 2, how IMBBH sources of different representative total masses ($10^{3}\,M_{\odot}$, $5\times 10^{3}\,M_{\odot}$, $10^{4}\,M_{\odot}$) at a fixed luminosity distance of $3$ Gpc will evolve in the LISA band. Considering observation time $T_{\rm ob}=4\,{\rm year}$ till ISCO, these sources will inspiral in the LISA band starting from the frequencies corresponding to black markers, merge at frequencies corresponding to magenta markers.
Let $f_{\rm ref}$ be the reference frequency to which the binary has evolved, when the ST hair is implanted on the SMBH. We choose this frequency to be $4\,{\rm mHz}$, which is close to the high sensitivity regime of LISA. Indeed, the binaries are still $375$ days, $26$ days and $8$ days (shown by purple markers), respectively, away from reaching their respective ISCOs.
Figure 2: The evolution of the characteristic (dimensionless) strain Moore et al. (2015) (inclination-averaged) of three representative IMBBH sources with equal-mass components (solid lines) shown in comparison to the characteristic (dimensionless) strain corresponding to the noise PSD of LISA (dashed curve). Luminosity distance of the IMBBH sources is fixed at $3$ Gpc. The magenta markers denote the GW strain and GW frequency of the IMBBH sources at the corresponding ISCO. The purple markers at $f_{\rm ref}=4\,{\rm mHz}$ denote the time till ISCO. The black markers represent the characteristic strain, and GW frequency, $T_{\rm obs}=4$ years before ISCO.
IV Model selection
@@
[…]
@@
where ${\rm FF}$ is defined as the match between the two waveform models $\tilde{h}_{b}(f)$ ($\tilde{h}_{a}(f)$) under two different hypotheses $\mathcal{H}_{b}$ ($\mathcal{H}_{a}$) maximized over the intrinsic source parameters $\vec{\theta}$ and can be written as Apostolatos (1995); Vallisneri (2012):
$\text{FF}\equiv\langle\tilde{h}_{a}(f)|\tilde{h}_{b}(f)\rangle_{\text{max}(\vec{\theta})}\,.$ (26)
@@
[…]
@@ V Results
Figure 3: The ${\rm\log}({\rm ln\mathcal{B}}^{\rm b}_{\rm a})$ to determine whether the model of GWs with (b)/without (a) imprints of black hole memory is preferred, is shown in the color scale. These Bayes factors are evaluated on a grid of total binary mass (${\rm\log}(M_{\rm bin}/M_{\odot})$) and mass of supermassive black hole (${\rm\log}(M/M_{\odot})$). GW frequency corresponding to the time when supertranslation hair is imparted to SMBH is fixed at $f_{\rm ref}=4\,{\rm mHz}$. Supertranslation charge $q$ is varied across the different panels shown. This figure corresponds to the case when ST memory leads to an increase (corresponding to $\cos(2\theta)=-1$ ) in binary separation.
@@
Having introduced the waveform models in Sec. II.3, Bayesian inference in Sec. III and model comparison in Sec. IV, we discuss the results here. We construct a $21\times 21$ grid uniform-in-log of the total mass $M_{\rm{bin}}$ of IMBBH and uniform-in-log of the mass $M$ of SMBH. The masses of the IMBBH and SMBH are restricted to the ranges $10^{3}\mbox{--}10^{4}\,\rm{M_{\odot}}$ and $10^{6}\mbox{--}10^{9}\,\rm{M_{\odot}}$, respectively. We fix the mass-ratio at q $=1$ (corresponding to $m_{1}=m_{2}$). The binary components of the IMBBH source are assumed to be nonspinning. The luminosity distance $D_{L}$ of the IMBBH source (and SMBH as well) is chosen to be $3$ Gpc. We calculate the SNRs of the simulated sources within the LISA band and find that the SNRs fall in the range ($50-300$). These SNRs are high enough to validate the high-SNR approximation used in Eq. (25) of Sec. IV for simplifying the calculation of the Bayes factor Vallisneri (2008); Cokelaer (2008); Rodriguez et al. (2013).
@@
The supertranslation charge $q$ appearing in Eq. (11) is chosen from a set of discrete values: ($10^{-7}$, $10^{-6}$, $10^{-4}$, $10^{-3}$) and the reference GW frequency $f_{\rm{ref}}$, corresponding to the time at which ST memory is imparted to the binary, is chosen to be $4\,\rm{mHz}$. This corresponds to the bottom of the noise PSD “bucket” (Fig. 2) where the LISA detector is the most sensitive.
@@
To show that the GW signal prefers the model with ST memory over the model without ST memory, we calculate Bayes-factor for all the simulated sources using the parameter setup discussed. Figure 3 (corresponding to $\cos 2\theta=-1$ in Eq. (11)) shows the scatter-plot for Bayes-factor in the $M-M_{\rm{bin}}$ plane. The different panels in the figure pertain to different values of supertranslation charge as mentioned. We see that Log-Bayes-Factor increases with increasing total mass $M_{\rm{bin}}$ of the IMBBH which is due to an increase in the optimal SNR of the GW signal. Given a total mass $M_{\rm{bin}}$ of the binary, Log-Bayes-Factor shows weak dependence on mass $M$ of SMBH. Increasing the value ST charge $q$ increases the strength of supertranslation memory as shown in different panels of Fig. 3. When the ST charge $q$ is higher, we get higher Log-Bayes-Factor for given masses $M_{\rm{bin}}$, $M$ of the binary and SMBH, respectively.
For all chosen values of ST charge $q$, Log-Bayes-Factor values are in the range $10^{3}\mbox{--}10^{5}$ as shown in different panels of Fig. 3. This shows that the near-horizon ST memory can be confidently detected in the LISA band. We also computed Log-Bayes-Factor corresponding to $\cos 2\theta=1$ and sky-averaged value of $\cos 2\theta$ in Eq. (11). We do not find any significant difference (at order of magnitude) compared to $\cos 2\theta=-1$.
VI Conclusion

### v2

@@ III Frequency modulation of the binary
In the previous subsection, we argued that an inspiralling intermediate mass binary black hole (IMBBH) in the vicinity of horizon of supermassive black hole (SMBH) acts as a detector for detecting supertranslation (ST) memory. IMBBH loses energy and angular momentum to GWs which can be observed by LISA. The separation of the binary components $L=2a$ (we call it IMBBH separation) at any instant during inspiral can be related to GW frequency $f$ emitted at that time via Kepler’s law as Peters and Mathews (1963); Peters (1964):
@@
where $a$ is the radius of the binary, $1+z=(1+z_{c})(1+z_{\rm g})$ accounts for the cosmological and gravitational redshifts of the IMBBH, and $M_{\mathrm{bin}}$ is the source-frame total mass of the IMBBH. Note that the gravitational redshift $z_{g}$ of GWs from IMBBH placed near the ISCO of central Schwarzschild SMBH is approximately $0.225$. Whereas, the GWs from IMBBH in the vicinity of horizon/ISCO of an extremal Kerr SMBH can either be redshifted or blueshifted depending on the direction of emission of GWs. The net redshift (here denoted by $z_{\rm g}$) is restricted to the range [0.4,0.7][-0.4,0.7]. (See Igata et al. (2020) and Figure $6$ therein.)
Once the supertranslation ‘hair’ is implanted on the central SMBH, the binary separation in IMBBH changes as a result of ST memory as given by Eqs. ( 10, II.3). The sudden change in the IMBBH separation leads to a corresponding sudden change in the emitted GW frequency. This instantaneous frequency jump in the GW signal of the IMBBH should be observable. The detection of the ST memory can be quantified by calculating the mismatch between the original GW signal when there is no ST memory and the signal modulated due to the ST memory.
GWs in general relativity have two polarizations i.e., plus $h_{+}$ and cross $h_{\times}$. For quasicircular binaries, the evolution of these polarizations is determined by the intrinsic source parameters ($\vec{\theta}$) i.e., component masses ($m_{1}$, $m_{2}$) and spins ($\vec{\chi}_{1}$, $\vec{\chi}_{2}$); and the extrinsic source parameters ($\vec{\lambda}_{h}$) i.e., luminosity distance $D_{L}$ to the source, inclination angle $\iota$ of the binary with respect to the observer’s line of sight, time $t_{c}$ and phase $\phi_{c}$ at the coalescence. Note that instead of ($m_{1}$, $m_{2}$), it is more natural to use the chirpmass $\mathcal{M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ and mass ratio ${\bf q}=m_{2}/m_{1}$  [^5] Here q should not be confused with the supertranslation charge $q$ in Eqs. (II.2,10). in the parameter estimation as the shape of gravitational waveform is more sensitive to the chirpmass Cutler and Flanagan (1994). An interferometric GW detector, such as LISA, records the time-domain strain $h(t;\vec{\theta},\vec{\lambda})$ due to an incoming GW signal, which can be expressed as a linear combination of the plus ($h_{+}$) and cross ($h_{\times}$) polarizations of GWs weighted by the respective antenna responses ($F_{+},F_{\times}$):
@@
$h_{b}(t;\vec{\theta},\vec{\lambda})=h_{a}(t+\Delta t;\vec{\theta},\vec{\lambda})\,.$ (19)
@@
$\hat{\mathcal{A}}=\sqrt{{5\over 24}}\sqrt{{4\over 5}}{\sqrt{\eta}M_{\rm bin}^{5/6}(1+z)^{5/6}\over\pi^{2/3}D_{L}}\,.$ (21)
@@
where $t(f)$ is the time corresponding to GW frequency $f$, $f_{\rm in}$ and $f_{\rm fi}$ are, respectively, the GW frequencies at which binary is emitting before and after supertranslation hair is implanted on the central SMBH and ${\rm\mathcal{M}_{c}}=M_{\rm bin}\eta^{3/5}$ is the chirp mass of the binary. Eq. (III) can be expressed in terms of IMBBH mass $M_{\rm bin}$, initial IMBBH separation $L_{0}$ and upto the linear order in dimensionless supertranslation charge $q_{0}$ as
$\Delta t\propto{q_{0}\over f_{\rm ref}^{8/3}\eta(1+z_{\rm g})^{5/3}(1+z_{c})^{5/3}M_{\rm bin}^{5/3}}\,$, (25)
where $z_{\rm g}$ is the gravitational redshift of gravitational waves emitted by IMBBH in the neighbourhood of SMBH. Note that in order to simplify that calculations, we assume the frequency jump due to ST memory from $f_{\rm in}$ to $f_{\rm fi}$ to be discrete. This corresponds to the maximum time shift in the IMBBH GW signal throughout the shock wave pulse and henceforth, the Bayes Factors obtained with this assumption will be considered as the upper limit.
IV Bayesian Parameter Inference
@@
$p(d)=\int\Pi(\vec{\theta},\vec{\lambda})p(d|\vec{\theta},\vec{\lambda})d\vec{\theta}d\vec{\lambda}\,.$ (27)
Assuming the detector noise to be stationary and Gaussian, the likelihood can be written as Cutler and Flanagan (1994):
@@
where $d$ is the strain data of the detector and $\langle.|.\rangle$ is the noise-weighted inner product defined as:
$\langle A|B\rangle=2\int^{\rm{f_{high}}}_{\rm{f_{low}}}{A(f)B^{*}(f)+A^{*}(f)B(f)\over S_{n}(f)}df\,.$ (28)
@@
$\rho=\sqrt{4\int^{\infty}_{0}{\left|\tilde{h}(f)\right|^{2}\over S_{n}(f)}df}\,.$ (29)
@@
The lower and upper cutoff frequencies ($f_{\rm low}$ and $f_{\rm high}$) used in calculating the integral in Eq. (28) are determined by the detector sensitivity, observation time, and the source properties. The $f_{\rm low}$ is defined as $f_{\rm low}={\rm max}[10^{-4},f_{\rm year}]\,{\rm Hz}$, where $f_{\rm year}$ is the GW frequency $\rm T_{obs}$} years before the binary emits at the frequency corresponding to the innermost stable circular orbit (ISCO). This frequency, in the quadrupole radiation approximation, is determined by the chirp mass $\mathcal{M}_{c}$} and the observation time prior to ISCO $T_{\rm obs}$ as Berti et al. (2005):
@@
We assume $T_{\rm obs}=4$ year in our calculation. The upper cut-off frequency is given by $f_{\rm high}={\rm min}[0.1,f_{\rm ISCO}]\,{\rm Hz}$, where $f_{\rm ISCO}=1/({6^{3/2}\pi(1+z)M_{\rm bin}})$ is the GW frequency corresponding to ISCO of a test particle orbiting around a Schwarzschild black hole of mass $M_{\rm bin}$.
We have displayed in Fig. 2, how IMBBH sources of different representative total masses ($10^{3}\,M_{\odot}$, $5\times 10^{3}\,M_{\odot}$, $10^{4}\,M_{\odot}$) at a fixed luminosity distance of $3$ Gpc will evolve in the LISA band. Considering observation time $T_{\rm ob}=4\,{\rm year}$ till ISCO, these sources will inspiral in the LISA band starting from the frequencies corresponding to black markers, merge at frequencies corresponding to red markers.
Let $f_{\rm ref}$ be the reference frequency to which the binary has evolved, when the ST hair is implanted on the SMBH. We choose this frequency (say, for example, $4\,{\rm mHz}$) within the higher sensitivity regime of LISA band. Indeed, the binaries are still $375$ days, $26$ days and $8$ days (shown by blue markers), respectively, away from reaching their respective ISCOs.
Figure 2: The evolution of the characteristic (dimensionless) strain Moore et al. (2015) (inclination-averaged) of three representative IMBBH sources with equal-mass components (solid lines) shown in comparison to the characteristic (dimensionless) strain corresponding to the noise PSD of LISA (dashed curve). Luminosity distance of the IMBBH sources is fixed at 33 Gpc. The red markers denote the GW strain and GW frequency of the IMBBH sources at the corresponding ISCO. The blue markers at $f_{\rm ref}=4\,{\rm mHz}$ denote the time till ISCO. The black markers represent the characteristic strain, and GW frequency, $T_{\rm obs}=4$ years before ISCO.
V Model selection
@@
$\mathcal{B}^{b}_{a}={p(d|\mathcal{H}_{b})\over p(d|\mathcal{H}_{a})}={\int p(\vec{\theta}|\mathcal{H}_{b})p(d|\vec{\theta},\mathcal{H}_{b})d\vec{\theta}\over\int p(\vec{\theta}|\mathcal{H}_{a})p(d|\vec{\theta},\mathcal{H}_{a})d\vec{\theta}}\,.$ (31)
@@
where $\rm FF$} is defined as the match $\mathcal{M}$ between the two waveform models $\tilde{h}_{b}(f)$ ($\tilde{h}_{a}(f)$) under two different hypotheses $\mathcal{H}_{b}$ ($\mathcal{H}_{a}$) maximized over the intrinsic source parameters $\vec{\theta}$ and can be written as Apostolatos (1995); Vallisneri (2012):
$\text{FF}\equiv\mathcal{M}_{\text{max}(\vec{\theta})}\,.$ (33)
@@
$\langle\tilde{h}_{a}(f)|\tilde{h}_{b}(f)\rangle$ $=4\int^{f_{\rm ref}}_{f_{\rm low}}{\left|\tilde{h}_{a}(f)\right|^{2}\over S_{n}(f)}df$ (36)
$+2\int^{f_{\rm high}}_{f_{\rm ref}}{2\left|\tilde{h}_{a}(f)\right|^{2}\cos({2\pi f\Delta t})\over S_{n}(f)}df$
@@ VI Results
Figure 3: The Log-Bayes-Factor $({\rm ln\mathcal{B}}^{\rm b}_{\rm a})$ to determine whether the model of GWs with (b)/without (a) imprints of black hole memory for Schwarzschild black hole is preferred, is shown in the color scale. These Bayes factors are evaluated on a grid of total binary mass ($M_{\rm bin}$) and mass of supermassive black hole ($M$). GW frequency corresponding to the time when supertranslation hair is imparted to SMBH is fixed at $f_{\rm ref}=3\,{\rm mHz}$ (top row) or $f_{\rm ref}=8\,{\rm mHz}$ (bottom row). Dimensionless supertranslation charge $q_{0}$ is varied from left ($q_{0}=10^{-8}$) to right ($q_{0}=10^{-5}$) panels as shown. This figure corresponds to the case when ST memory leads to an increase (corresponding to $\cos(2\theta)=-1$ in Eq. (10)) in binary separation.
Figure 4: Same as Figure 3 but for extremal Kerr black hole (redshift $z_{g}=-0.4$) and this Figure corresponds to the case when ST memory leads to an increase (corresponding to $\cos(2\theta)=0$ and $\cos(\phi)=1$ in Eq. (II.3) ) in binary separation.
Figure 5: Same as Figure 3 but for extremal Kerr black hole (redshift $z_{g}=0.7$) and this Figure corresponds to the case when ST memory leads to an increase (corresponding to $\cos(2\theta)=0$ and $\cos(\phi)=1$ in Eq. (II.3) ) in binary separation.
@@
Having introduced the waveform models in Sec. III, Bayesian inference in Sec. IV and model comparison in Sec. V, we discuss the results here. We construct a $21\times 21$ grid uniform-in-log of the total mass $M_{\rm{bin}}$ of IMBBH and uniform-in-log of the mass $M$ of SMBH. The masses of the IMBBH and SMBH are restricted to the ranges $10^{3}\mbox{--}10^{4}\,\rm{M_{\odot}}$ and $10^{7}\mbox{--}10^{9}\,\rm{M_{\odot}}$, respectively. We fix the mass-ratio at q =1=1 (corresponding to $m_{1}=m_{2}$). The binary components of the IMBBH source are assumed to be nonspinning. The luminosity distance $D_{L}$ of the IMBBH source (and SMBH as well) is chosen to be $3$ Gpc. We calculate the SNRs of the simulated sources within the LISA band and find that the SNRs fall in the ranges (5040050-400), (3022030-220) and (7552075-520) for Swarzschild case, extreme Kerr cases with redshifts $z_{\rm g}=-0.4$ and $z_{\rm g}=0.7$, respectively. These SNRs are high enough to validate the high-SNR approximation used in Eq. (32) of Sec. V for simplifying the calculation of the Bayes factor Vallisneri (2008); Cokelaer (2008); Rodriguez et al. (2013).
@@
The supertranslation charge $q_{0}$ appearing in Eqs. (10,II.3) is chosen from a representative set of discrete values: ($10^{-8}$, $10^{-5}$) and the reference GW frequency $f_{\rm{ref}}$, corresponding to the time at which ST memory is imparted to the binary, is chosen to be ($3\,\rm{mHz}$, $8\,\rm{mHz}$). This corresponds to the significant fraction of bottom of the noise PSD “bucket” (Fig. 2) where the LISA detector is the most sensitive.
@@
To show that the GW signal prefers the model with ST memory over the model without ST memory, we calculate Bayes-factor for all the simulated sources using the parameter setup discussed. Figure 3 (corresponding to $\cos 2\theta=-1$ in Eq. (10)) shows the scatter-plot for Log-Bayes-Factor in the $M-M_{\rm{bin}}$ plane. The different panels in each row of Fig. 3 pertain to different values of dimensionless supertranslation charge $q_{0}$ while keeping the reference frequency fixed as mentioned. We see that Log-Bayes-Factor shows a non-trivial trend with the total mass $M_{\rm{bin}}$ of the IMBBH. This is because Log-Bayes-Factor as given in Eq. (32) depends on two factors, fitting factor $FF$ [see Eqs. ( 34- 36)] and square of the SNR $\rho^{2}$. $FF$ has a sinusoidal dependence on time-shift $\Delta t$ with $\Delta t$ having a decreasing trend with increase in the binary mass $M_{\rm bin}$ [see Eq. (25)] and SNR $\rho$ increases with increase in mass of the binary $M_{\rm bin}$. These two competing factors decide the net dependence of Log-Bayes-Factor on the binary mass $M_{\rm bin}$ which also depends on the combination of other parameters like dimensionless ST charge $q_{0}$, reference frequency $f_{\rm ref}$, gravitational/relativistic redshift/blueshift $z_{\rm g}$ and symmetric mass ratio $\eta$. Given a total mass $M_{\rm{bin}}$ of the binary, Log-Bayes-Factor doesn’t shows any dependence on mass $M$ of SMBH. This is because $\delta\phi$ and $\delta\theta$ in Eqs. (10,II.3) are inversely proportional to $M^{2}$ making $\Delta L$ and hence $\Delta t$ [Eq. (25)] independent of M. Increasing the value of ST charge $q_{0}$ increases the strength of supertranslation memory as shown in different panels of each row of Fig. 3. When the ST charge $q_{0}$ is higher, we get higher Log-Bayes-Factor for given masses $_{\rm{bin}}, MM$ of the binary and SMBH, respectively, as $\Delta L\propto q_{0}$. Moreover, as we move from top row ($f_{\rm ref}=3\,{\rm mHz}$) to bottom row ($f_{\rm ref}=8\,{\rm mHz}$) of Fig. 3, Log-Bayes-Factor reduces by nearly two orders of magnitude. This is because $\delta\phi$ and $\delta\theta$ in Eq. (10) are directly proportional to $L_{0}^{2}$ and hence making $\Delta L$ proportional to $L_{0}$ ($L_{0}\propto f_{\rm ref}^{-2/3}$).
Figures (4, 5) correspond to the case when IMBBH is nearly at the ISCO/horizon ($r\approx\,(1+0.001)M$) of an extremal Kerr SMBH with extreme redshift choices of $z_{\rm g}=-0.4$ and $z_{\rm g}=0.7$, respectively, from the allowed redshift range $z_{\rm g}\in[-0.4,0.7]$ Igata et al. (2020). The trends in the Log-Bayes-Factor with respect to dimensionless ST charge $q_{0}$, reference frequency $f_{\rm ref}$, IMBBH mass $M_{\rm bin}$ and SMBH mass $M$ hold the similar explanations as for Fig. 3. Comparing the Log-Bayes-Factor in Fig. 4 with the correspondings numbers in Fig. 3, there is nearly three(one) order(s) of improvement across the chosen values of reference frequency $f_{\rm ref}$ with fixed $q_{0}=10^{-8}$ ($10^{-5}$). Similarly, comparing the Log-Bayes-Factor in Fig. 5 with the corresponding numbers in Fig. 3, there is approximately two(one) orders(s) of improvement across the chosen values of reference frequency $f_{\rm ref}$ with fixed $q_{0}=10^{-8}$ ($10^{-5}$). This suggests that the detection prospects of black hole memory in the vicinity of the horizon of extremal Kerr SMBH are better compared to corresponding detection probability near ISCO of Schwarzschild SMBH. Moreover, it is important to note that we also computed the Log-Bayes-Factor for the Kerr Case with intermediate values of redshift $z_{\rm g}=0.2,0.5$ and we didn’t notice any drastic change in the results when comapred to the Kerr case with redshift $z_{\rm g}=0.7$ [corresponding to Fig. 5].
For all chosen values of dimensionless ST charge $q_{0}$ and reference frequency $f_{\rm ref}$, Log-Bayes-Factor is the best for a larger supertranslation charge ($q_{0}=10^{-5}$) and a smaller reference frequency ($f_{\rm ref}=3\,{\rm mHz}$) for both the Schwarzschild ($\ln\mathcal{B}^{b}_{a}\approx 10^{3}-10^{4}$) and extremal Kerr ($\ln\mathcal{B}^{b}_{a}\approx 10^{4}-10^{5}$) SMBHs. This is because $\Delta t$ in Eq. (25) is directly proportional to $q_{0}$ and the dependence on reference frequency $f_{\rm ref}$ is demonstrated via $\Delta t\propto f^{-8/3}_{\rm ref}$. The large Log-Bayes-Factor obtained for a significant portion of the parameter space considered, for both Schwarzchild and extremal Kerr SMBHs, show that the near-horizon ST memory can be confidently detected in the LISA band.
VII Conclusion